OpenStudy (anonymous):

The sum of two consecutive intergers is 153. Which equation can be used to find the first interger n?

OpenStudy (anonymous):

A. 2n + 1 = 153 B. 2n = 153 C. n + 1 =153 D. 2n + 2 = 153

OpenStudy (raden):

let the first number is n, so the 2nd is n+1 add of them, what u get ....

OpenStudy (zehanz):

If the first is n, the consecutive is n+1. Adding them gives 153. Try to make an equation out of this...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait if what first number is n

OpenStudy (anonymous):

so the answer to this question is a beacause there must b a # in front of n and n is supposed to be added by 1 to equal 153 so am i right

OpenStudy (zehanz):

You have found the right answer: n+n+1 = 153 2n+1=153 Answer A.

OpenStudy (anonymous):

ya i feel smart i am definently getting the galaxy s3

OpenStudy (zehanz):

;)